Each of two road graders can pave a certain section of the road in 15 hours working alone. After two machines have been working for 2 hours they are joined by an older model that would take 20 hours to pave the section road by itself. How long after the first two machines began working will it take the three machines to pave the section of the road?
Answer by stanbon(75887)

Each of two road graders can pave a certain section of the road in 15 hours working alone. After two machines have been working for 2 hours they are joined by an older model that would take 20 hours to pave the section road by itself. How long after the first two machines began working will it take the three machines to pave the section of the road?
------------------------
Rate of each road grader = 1/15 job/hr
-----
Rate for 2 of the road graders working together = 2/15 job/hr
---------------------------
Older model rate = 1/20 job/hr
--------------------------------------
How long after the first two machines began working for 2 hrs will it take the three machines to pave the section of the road?
---
Let the # of hrs required be "x":
2(2/15) + x(1/15 + 1/15 + 1/20) = 1 job
-------
(4/15) + x[(4 + 4 + 3)/60] = 1
---
(11/60)x = 11/15
x = 4 hrs. (Additional time needed to finish the job)
===================================
Cheers,
Stan H.
============